OpenStudy (anonymous):

PLEASE HELP Write the equation in vertex form? (working with parabolas). 21.) vertex (-3,7) point (-2,-5) 22.) vertex (4,0) point (-6,-3) 23.) vertex (-2,-5) point (-6,0) 24.) vertex (1,3) point (2,5)

OpenStudy (anonymous):

Lets see the equation for Standard form is. \[y = a(x-h)^{2} + k, \\ \text{Your h is the x coordinate, k is the y coordinate (of the vertex)}\\ \text{plug in your points given for 21. x = -2 , y =-5, h =-3, k = 7. }\\ \text{enjoy!}\]

OpenStudy (anonymous):

that is actually vertex form

OpenStudy (anonymous):

hold on im thinking.

OpenStudy (anonymous):

20.) OPEN-ENDED write a quadratic function that has a maximum value?

OpenStudy (anonymous):

vertex is (-3,7) plug that into y=a(x-h)^2+k you get: y=a(x-(-3))^2+7 now time to find a. plug the rest in. (-2-5) -5=a((-2)-(-3))^2+7 Simplify. -5=a(-2+3)^2+7 -5=a(1)^2+7 -5=8a Divide both sides by 8 -5/8=a put into original equation y=-5/8(x+3)^2+7 pretty sure thats right. not 100%

OpenStudy (anonymous):

thanks I have 2 last questions numbers 11,12 then im done.. Thanks!! (=

OpenStudy (anonymous):

np! do you mind taking a look at my last question? maybe you read the poetry im tested on.